### v1

# Escaping the BLEU Trap: A Signal-Grounded Framework with Decoupled Semantic Guidance for EEG-to-Text Decoding

Abstract
Decoding natural language from non-invasive EEG signals is a promising yet challenging task. However, current state-of-the-art models remain constrained by three fundamental limitations: Semantic Bias (mode collapse into generic templates), Signal Neglect (hallucination based on linguistic priors rather than neural inputs), and the BLEU Trap, where evaluation metrics are artificially inflated by high-frequency stopwords, masking a lack of true semantic fidelity. To address these challenges, we propose SemKey, a novel multi-stage framework that enforces signal-grounded generation through four decoupled semantic objectives: sentiment, topic, length, and surprisal. We redesign the interaction between the neural encoder and the Large Language Model (LLM) by injecting semantic prompts as Queries and EEG embeddings as Key-Value pairs, strictly forcing the model to attend to neural inputs. Furthermore, we move beyond standard translation metrics by adopting N-way Retrieval Accuracy and Fréchet Distance to rigorously assess diversity and alignment. Extensive experiments demonstrate that our approach effectively eliminates hallucinations on noise inputs and achieves SOTA performance on these robust protocols. Code will be released upon acceptance at https://github.com/xmed-lab/SemKey.
1 Introduction
Figure 1: Top: Previous models (Liu et al., 2025) first exhibit semantic bias, overfitting to generic templates (e.g., “He was a…” ) to artificially inflate BLEU scores despite severe semantic misalignment (the “BLEU Trap”). Furthermore, the noise test (Row 3) exposes signal neglect: the model continues to hallucinate fluent text even from pure Gaussian noise, proving that generation is driven by linguistic priors rather than the input signal. Bottom: In contrast, SemKey(Our Method) demonstrates strict signal dependency. It generates diverse, semantically faithful text for valid inputs while correctly yielding disordered tokens for noise, verifying that decoding is genuinely grounded in neural signals.
Decoding natural language from neural activity is a pivotal frontier in Brain-Computer Interface (BCI) research, offering a lifeline for individuals with aphasia or locked-in syndrome to restore communication (Défossez et al., 2023; Willett et al., 2023). While invasive modalities like Electrocorticography (ECoG) and Stereoelectroencephalography (sEEG) demonstrate high fidelity (Metzger et al., 2023; Feng et al., 2025), their invasive procedures make them unsuitable for widespread use in general-purpose BCI applications. Consequently, non-invasive Electroencephalography (EEG) remains the most promising modality for widespread adoption due to its portability and safety.
EEG-to-text decoding (Kostas et al., 2021; Wang & Ji, 2022; Duan et al., 2023; Liu et al., 2024), which aims to translate neural signals into natural language descriptions, holds great promise for enabling direct brain–computer communication and assisting individuals with severe speech or motor impairments, and has attracted growing attention. However, this field currently faces a validity crisis: recent work (Jo et al., 2024) reveals that reported successes are largely artifacts of Teacher Forcing, i.e., they depend on ground-truth tokens to stabilize generation and cannot produce coherent sequences under fully autoregressive decoding, rendering prior methods impractical. While GLIM (Liu et al., 2025) achieves state-of-the-art open-vocabulary performance by eliminating teacher forcing, our analysis reveals three persistent limitations. First, semantic bias and repetition: models fail to capture distinct semantics, instead overfitting to generic templates (e.g., “He was…”) and exhibiting recursive loops, as shown in Figure 1 and Appendix Table g&LABEL:tab:full_qualitative_comparison. Second, signal neglect: despite GLIM’s improvements, our noise perturbation experiments reveal a continued heavy reliance on linguistic priors rather than true neural decoding (see Figure 4 and Table h). Finally, the “BLEU Trap”: current evaluation protocols relying on translation metrics are flawed. Models can achieve competitive scores simply by matching stopwords and templates, artificially masking poor semantic fidelity (Callison-Burch et al., 2006; Liu et al., 2016) (Table g).
To this end, we propose SemKey, a novel two-stage framework enforcing semantic guidance and strict signal grounding. In the first stage, the Parallel Multi-task Attribute Extraction stage is inspired by neuroscientific findings that high-level semantics form robust representations (Lerner et al., 2011). Therefore, we decouple the decoding process by explicitly predicting four attributes (sentiment, topic, length, and surprisal) to mitigate semantic bias and reduce repetitive templates. These serve as stable semantic anchors to steer the LLM, effectively breaking the cycle of bias and repetition. Then, in the second stage, the Multi-Perspective Active Retrieval Decoding stage, to mitigate signal neglect, we redesign the encoder-LLM interaction via a Q-K-V Injection mechanism. Recognizing that standard cross-attention often fails to override the LLM’s strong language modeling objective, we treat EEG embeddings as Keys/Values explicitly queried by multi-perspective semantic prompts obtained from the first stage. This mechanism transforms decoding from passive conditioning to active information retrieval, ensuring generation remains strictly grounded in neural input. Furthermore, based on our observation that previous baselines achieve high BLEU scores primarily through template memorization rather than genuine decoding, we advocate moving beyond this “BLEU trap”. We establish a comprehensive evaluation protocol using rigorous Retrieval and Distribution metrics. Experiments demonstrate that SemKey achieves state-of-the-art performance, outperforming the GLIM baseline by 42.7% in Self-BLEU, 54.4% in FD, and 36.4% in 24-Way Retrieval Accuracy.
2 Related Work
EEG-to-Text Decoding: Evolution and Frontiers. Early approaches relied on handcrafted features like Mel-frequency cepstral coefficients (MFCCs) (Cooney et al., 2018) and wavelet transforms (Panachakel et al., 2020), necessitating extensive domain expertise. The advent of deep learning introduced automated feature extraction via CNNs and LSTMs (Zhang et al., 2018) to capture spatio-temporal dynamics. A paradigm shift occurred with Transformer-based pipelines: (Wang & Ji, 2022) proposed projecting EEG signals into a pre-trained BART latent space, and building on this, DeWave (Duan et al., 2023) introduced an end-to-end framework processing raw brainwaves for direct translation.
More recently, the integration of Large Language Models has spurred new directions. Thought2Text (Mishra et al., 2025) explores leveraging LLMs for EEG encoding, while WaveMind (Zeng et al., 2025) proposes a conversational foundation model aligned across textual and visual modalities. However, our work differs from these distinct streams: unlike WaveMind, which relies on image stimuli that offer richer and more direct semantic grounding, we tackle the different task of pure EEG-to-Text decoding from abstract textual stimuli without such visual aids.
The Validity Crisis and Methodological Gap. Notwithstanding architectural advancements, the field faces a crisis of validity. The pivotal study “Are EEG-to-Text Models Working” (Jo et al., 2024) exposed that high BLEU scores in prior arts are often artifacts of Teacher Forcing, where models ignore the noisy EEG encoder in favor of the LLM’s linguistic priors. While GLIM (Liu et al., 2025) recently addressed this by establishing the first effective pipeline without teacher forcing, it—along with its predecessors—still suffers from “signal neglect,” leading to semantic collapse in open-ended generation (as discussed in Section 1).
To bridge this gap between open-ended generation and faithful neural decoding, SemKey departs from the prevailing reliance on implicit latent alignment. We propose a Coarse-to-Fine decoding paradigm to enforce validity: First, we introduce an explicit, attribute-driven mechanism where high-level semantic anchors (e.g., Sentiment, Topic) actively steer the decoding trajectory. This serves as a hard constraint to prevent semantic collapse, ensuring adherence to the global semantic intent. Complementing this high-level guidance, we reformulate the fine-grained LLM interaction via a Q-K-V Injection mechanism. By forcing EEG embeddings to act as Keys and Values, we enforce a strict token-level dependency on neural signals.
3 Method
Figure 2: The overall architecture of the SemKey framework following a ”Guidance-Generation” paradigm. Stage 1 (Attribute Extraction): The EEG encoder is optimized via multi-task learning to explicitly decouple high-level semantic attributes (Sentiment, Topic, Length, Surprisal) alongside standard alignment. Stage 2 (Generative Decoding): These predicted attributes structure a semantic prompt to guide generation. Crucially, the Q-K-V Injection mechanism enforces signal dependency by using the text prompt as the Query and projected EEG embeddings as Keys and Values.
SemKey is a two-stage framework designed to mitigate semantic bias and signal neglect. It comprises: (1) a Parallel Multi-Task Attribute Extraction stage that learns robust EEG representations aligned with the linguistic latent space while explicitly decoupling high-level semantic attributes; and (2) a Multi-Perspective Active Retrieval Decoding stage that actively exploits these representations to produce coherent and faithful text.
3.1 Stage 1: Parallel Multi-Task Attribute Extraction
Based on (Liu et al., 2025), we employ a multi-task learning strategy to explicitly decouple high-level semantic labels. The EEG encoder transforms signals into latent embeddings, inheriting alignment and reconstruction tasks via a frozen Flan-T5 model (Chung et al., 2024). Concurrently, we introduce four auxiliary decoding heads to predict semantic attributes: Sentiment, Topic, Length, and Surprisal. These predictions serve as a robust semantic skeleton to guide the subsequent generation. As shown in Figure 2, this stage is structured into three key components: a prior-guided encoding stream, a parallel multi-task learning framework, and a unified training objective.
Prior-Guided EEG Encoding Stream. To mitigate the severe inter-subject variability and domain shifts inherent in raw EEG signals (Jiang et al., 2024), we adopt the Prior Label Injection strategy utilized in GLIM (Liu et al., 2025). Let $\mathcal{X}_{eeg}\in\mathbb{R}^{T\times C}$ denote the input EEG signals. Following this protocol, discrete priors $\mathcal{L}_{prior}$ (including Subject ID, Dataset source, and Task type) are aggregated into a global context vector $\mathbf{e}_{prior}$. This vector is then employed to dynamically modulate the intermediate features within the hierarchical Conformer encoder (Gulati et al., 2020; Jiang et al., 2024), ensuring the feature extraction process remains robust against individual noise and distribution shifts. The implementation details can be found in Appendix B.
After that, an Aligner projects the encoded states into the shared semantic dimension $d_{model}$, yielding distinct dual representations to serve different decoding granularities: (1) Sequence Embeddings ($\mathbf{E}_{eeg}\in\mathbb{R}^{L\times d_{model}}$), which preserve fine-grained temporal details for word-level reconstruction; and (2) Global Embedding ($\mathbf{v}_{eeg}\in\mathbb{R}^{d_{model}}$), obtained via attention pooling (Lee et al., 2019), which summarizes the holistic semantic information for high-level attribute prediction. Following GLIM (Liu et al., 2025), we adopt a reconstruction loss $\mathcal{L}_{recon}$ (standard autoregressive cross-entropy) to enforce precise token generation from the sequence embeddings $\mathbf{E}_{eeg}$, alongside an alignment objective $\mathcal{L}_{align}$ (comprising Contrastive and Commitment losses) to synchronize the EEG embeddings (Contrastive is calculated with global embedding ($\mathbf{v}_{eeg}$) and Commitment is calculated with sequence embeddings ($\mathbf{E}_{eeg}$)) with the text latent space.
Parallel Multi-Task Learning Framework. To enforce semantic richness in $\mathbf{v}_{eeg}$ and avoid the semantic collapse observed in prior works, we introduce a parallel multi-task framework (Wang et al., 2025b). While we retain the standard contrastive alignment ($\mathcal{L}_{align}$) and token reconstruction ($\mathcal{L}_{Recon}$) tasks to ensure basic linguistic capability, our core innovation lies in the Decoupled Attributes Prediction. We attach four parallel heads (MLPs) directly to $\mathbf{v}_{eeg}$ to explicitly disentangle high-level semantic features:
(1) Classification Objectives (Sentiment & Topic): Two classifiers predict the discrete labels for Sentiment and Topic. We optimize these using the Cross-Entropy loss $\mathcal{L}_{stm/tpc}=-\sum_{c=1}^{N_{C}}y_{c}\log(\hat{y}_{c})$, where $N_{C}$ is the number of classes, $y_{c}$ is the one-hot ground truth, and $\hat{y}_{c}$ is the predicted probability.
(2) Regression Objectives (Length & Surprisal): Two regressors predict continuous attributes: sentence Length and Surprisal (information density). These are optimized via the Mean Square Error (MSE) loss $\mathcal{L}_{len/spr}=\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}$. The regression targets are dynamically normalized to ensure stable convergence. These auxiliary tasks force the encoder to encode specific semantic properties into the global embedding, preventing it from degenerating into a generic representation.
Overall Training Objective. The model is trained end-to-end with a weighted sum of the alignment, reconstruction, and multi-perspective attribute losses:
$\begin{split}\mathcal{L}_{Stage1}=\lambda_{align}\mathcal{L}_{align}+\lambda_{recon}\mathcal{L}_{Recon}\\ +\lambda_{cls}(\mathcal{L}_{stm}+\mathcal{L}_{tpc})+\lambda_{reg}(\mathcal{L}_{len}+\mathcal{L}_{spr})\end{split}$ (1)
where $\lambda$ are hyperparameters balancing the contribution of each component. This composite objective results in an EEG Encoder capable of producing semantically disentangled and attribute-aware embeddings, laying the foundation for the generative stage.
3.2 Stage2: Multi-Perspective Active Retrieval Decoding
Prediction heads from Stage 1 are frozen, and the aligned EEG features are injected into the LLM via a learnable projector. To ensure fidelity, we introduce two key mechanisms: (1) Prompt Guidance, where the predicted attributes are structured into a natural language constraint; and (2) Q-K-V Injection, which re-routes the attention mechanism: the text prompt serves as the Query, interacting strictly with EEG embeddings provided as Keys and Values. This architecture transforms passive conditioning into an active retrieval process, compelling the LLM to actively query neural signals for every token generated. This stage follows an end-to-end paradigm, meaning that the EEG Encoder keeps trainable while the LLM Decoder keeps frozen. To address the “signal neglect” and “semantic bias” issue, we introduce a Guidance-Generation paradigm: a lightweight projector is imported, with a novel Q-K-V injection mechanism.
Multi-Perspective Prompt. To explicitly guide text reconstruction using the high-level semantics obtained in Stage 1, we leverage the decoupled attribute predictions to construct a structured natural language prompt, serving as a “semantic anchor” for the LLM (Keskar et al., 2019; Lester et al., 2021). Let $\mathcal{H}=\{\hat{y}_{len},\hat{y}_{tpc},\hat{y}_{stm},\hat{y}_{spr}\}$ be the set of predicted high-level attributes from Stage 1. We define a template function $\mathcal{T}(\cdot)$ that maps these attributes into a textual instruction sequence $\mathbf{X}_{prompt}$ as follows. System: Based on the following EEG signals, reconstruct the text. The length of the sentence is $\hat{y}_{len}$] words. The average surprisal value is $\hat{y}_{spr}$]. Sentiment: $\hat{y}_{stm}$]. Topic: $\hat{y}_{tpc}$]. This prompt $\mathbf{X}_{prompt}$ is tokenized and prepended to the decoder input as a prefix. In the context of the attention mechanism, the hidden states derived from this prompt act as the Queries ($Q$), actively seeking relevant information to fulfill the semantic constraints defined by the high-level attributes.
Q-K-V Injection Mechanism. Unlike prior approaches that simply concatenate brain embeddings with text tokens—often leading to signal neglect via self-attention (Wu et al., 2022; Mishra et al., 2025)—we employ a Q-K-V Injection strategy to strictly separate the information flow (Li et al., 2023; Alayrac et al., 2022). Let $\Phi_{self}(\cdot)$ denote the stacked causal self-attention layers of the frozen LLM decoder. During the autoregressive generation step $t$, the input sequence consists of the fixed structured prompt $\mathbf{X}_{prompt}$ (as a prefix) concatenated with the previously generated tokens $\mathbf{y}_{<t}$. The decoder first processes this sequence to obtain the contextualized hidden states $\mathbf{H}_{text}$:
$\mathbf{H}_{text}=\Phi_{self}([\mathbf{X}_{prompt};\mathbf{y}_{<t}])$ (2)
where $\cdot$$\cdot$] denotes the concatenation operation along the sequence dimension. These hidden states $\mathbf{H}_{text}$ then serve as the Queries for the subsequent cross-attention layers, specifically seeking information from the neural signals.
Let $\mathbf{E}_{eeg}\in\mathbb{R}^{L\times d_{model}}$ denote the projected EEG embeddings. The cross-attention mechanism is redefined as:
$Q$ $=\mathbf{H}_{text}\mathbf{W}_{Q},$ (3)
$K$ $=([\mathbf{v};\mathbf{E}_{eeg}]\mathbf{W}_{proj})\mathbf{W}_{K},$
$V$ $=([\mathbf{v};\mathbf{E}_{eeg}]\mathbf{W}_{proj})\mathbf{W}_{V}$
Mechanism Rationale: The reason for strictly assigning $\{\mathbf{v},\mathbf{E}_{eeg}\}$ as Keys and Values (interacting with linguistic Queries) is to establish an “Active Retrieval” paradigm (Bahdanau et al., 2014). In autoregressive generation, the linguistic context ($\mathbf{H}_{text}$) dictates what information is needed next (acting as the active Query), while the neural signals provide the content pool (serving as the passive Source) (Vaswani et al., 2017). By structuring $\{\mathbf{v},\mathbf{E}_{eeg}\}$ as the Key-Value memory bank, we incentivize the LLM to dynamically “lookup” and attend to specific neural segments based on current semantic requirements. Compared with simple concatenation strategies, this structural constraint mitigates the risk of input decoupling, as the decoder is architecturally encouraged to anchor its generation on the EEG-derived representations rather than relying solely on its internal language priors.
Data Augmentation. To further enhance the semantic robustness of the generative decoding, we adopt the Multiple Text Variants (MTV) strategy following (Liu et al., 2025). During the training of Stage 2, we utilize diverse paraphrases of the ground-truth text as augmentation targets, as shown in Figure 2. This approach encourages the model to capture invariant semantic information rather than overfitting to specific syntactic structures.
3.3 Escaping the “BLEU Trap”: A Holistic Evaluation Protocol
Standard evaluations in EEG-to-Text decoding heavily rely on n-gram matching metrics such as BLEU. However, our experiment shows that these metrics can be misleading in this domain. A model generating generic sentences can achieve high BLEU scores without extracting any real semantic information from the EEG, refer to Section 2. To avoid this “BLEU trap” and strictly evaluate both semantic alignment and generative diversity, we establish a comprehensive evaluation protocol consisting of three distinct categories:
Retrieval-based Alignment Metrics ($N$-Way Accuracy). To measure the pure semantic alignment capability of the EEG Encoder, we employ $N$-way Retrieval Accuracy, a standard protocol for evaluating modality alignment (Radford et al., 2021; Li et al., 2024). This metric evaluates whether the predicted EEG embedding can successfully retrieve the correct text from a pool of distractors. Specifically, we utilize the sentence transformers (Reimers & Gurevych, 2019) library to encode the ground-truth text and N1N-1 randomly selected negative samples into dense vectors. We then calculate the Cosine Similarity between the aligned EEG embeddings and these candidate text vectors.
To mitigate the variance from the random selection of negative samples, we conducted multiple experimental runs and reported the average accuracy to ensure stability. We report accuracy at four difficulty levels: 2-Way, 4-Way, 10-Way, and 24-Way Accuracy.
$\begin{split}\text{Acc}_{N}=\frac{1}{M}\sum_{i=1}^{M}\mathbb{I}\Big(&\text{CosSim}(\mathbf{e}_{eeg}^{(i)},\mathbf{v}_{text}^{(i)})\\ &>\max_{j\in\mathcal{N}_{neg}}\text{CosSim}(\mathbf{e}_{eeg}^{(i)},\mathbf{v}_{text}^{(j)})\Big)\end{split}$ (4)
where $M$ is the test set size, and $\mathcal{N}_{neg}$ is the set of N1N-1 negative samples. This metric strictly evaluates whether the model can distinguish specific semantic content from noise.
Semantic Substance Metric (Content Recall). High BLEU scores can be achieved by matching stopwords, while missing the core meaning. To evaluate the recovery of important information, we introduce Content Recall. We first filter out stopwords from both the reference and generated text, retaining only content-bearing words (nouns, verbs, adjectives). Content Recall calculates the percentage of ground-truth content words that successfully appear in the generated hypothesis. This metric penalizes “fluent but empty” hallucinations and rewards semantic fidelity (Ji et al., 2023; Maynez et al., 2020).
Generative Diversity & Quality Metrics. A major failure mode of previous EEG-to-Text models is semantic bias, where the model repeats a few safe sentences. To quantify the richness and non-repetitiveness of our generation, we utilize four key metrics.
First, to evaluate vocabulary usage and uniqueness, we employ Dist-1 & Dist-2 and Self-BLEU. Dist-$n$ metrics calculate the ratio of distinct unigrams and bigrams to the total number of generated tokens (Li et al., 2016), where higher values indicate broader vocabulary usage. Complementarily, Self-BLEU (Zhu et al., 2018) measures the BLEU score of each generated sentence against all others in the test set; a lower score indicates that the model generates unique samples for different EEG inputs rather than regurgitating memorized templates. Additionally, to assess structural diversity, we compute Head Entropy (Prefix Diversity). Following (Holtzman et al., 2019), we calculate the Shannon entropy of the initial bigrams (first two tokens) across all generated sentences. A higher Head Entropy signifies greater diversity in sentence initiations, preventing the model from collapsing into fixed opening templates. Finally, to evaluate semantic realism at the distribution level, we compute the Fréchet Distance (FD) (Semeniuta et al., 2018). We map all sentences to a high-dimensional semantic space using a pre-trained Sentence Transformer and calculate the distance between the multivariate Gaussian distributions of the generated ($\mu_{g},\Sigma_{g}$) and reference ($\mu_{r},\Sigma_{r}$) embeddings:
[…]
A lower FD indicates that the distribution of the decoded text closely matches the semantic manifold of the real text.
4 Experiments & Results
We conduct our experiments on the ZuCo 1.0 & ZuCo 2.0 benchmarks (Hollenstein et al., 2018, 2020), which serve as the canonical standard for EEG-to-Text decoding and are widely adopted by state-of-the-art studies. For comprehensive details regarding dataset splits, preprocessing protocols, and implementation specifics (including optimizer settings and hyperparameters), please refer to Appendix B.
4.1 Main Performance Comparison
Table 1: Main performance comparison on the ZuCo test set. We report retrieval accuracy, content recall, diversity (Dist-$n$, Head Entropy), and quality metrics. Best results are highlighted in bold. Note on Baselines: Results for GLIM (Liu et al., 2025) are obtained using their official public checkpoint. In contrast, results for EEG-to-Text (Wang & Ji, 2022) are marked with *, indicating they were reproduced using the authors’ code under the strict non-teacher-forcing setting, as originally reported metrics rely on invalid validation schemes (see Section 1). The last row shows the relative improvement (%) of our model compared to the GLIM baseline.
Retrieval Accuracy Substance Diversity Quality ($\downarrow$)
Model 2-Way 4-Way 10-Way 24-Way C. Recall Dist-1 Dist-2 H. Ent S-BLEU FD
EEG-to-Text* 74.8% 43.6% 19.0% 8.1% 2.1% 0.1% 0.1% 0.00 100.0% 1.65
GLIM 83.3% 57.5% 29.7% 14.3% 1.2% 3.0% 8.1% 2.18 90.8% 0.57
SemKey(Ours) 86.1% 62.8% 36.7% 19.5% 2.6% 5.3% 22.6% 6.21 52.0% 0.26
Rel. Improv. +3.4% +9.2% +23.6% +36.4% +116.7% +76.7% +179.0% +184.9% +42.7% +54.4%
(vs. GLIM)
Figure 3: Holistic Performance Evaluation. The Radar Chart illustrates the relative performance of our model (Orange) versus GLIM (Blue) and EEG-to-Text (Green). SemKey demonstrates a balanced superiority across all alignment, diversity, and content quality metrics.
We compare SemKey against two representative baselines: EEG-to-Text and GLIM. Notably, while DeWave is a prominent prior work, we exclude it from this comparison for two reasons: First, as discussed in (Jo et al., 2024), DeWave shares a nearly identical methodology with EEG-to-Text when the teacher forcing mechanism is removed during inference. Second, the source code has been removed from the official GitHub repository, where the authors explicitly acknowledged the issues of teacher forcing and signal neglect. Therefore, we adopt EEG-to-Text as the representative baseline for contrastive-based approaches before (Jo et al., 2024).
Quantitative Results. Table 1 and Figure 3 present the comprehensive performance comparison. SemKey demonstrates robust improvements across three key dimensions.
First, regarding Alignment & Semantic Substance, SemKey consistently outperforms baselines across all retrieval settings ($N=2,4,10,24$). Notably, the performance advantage widens as task difficulty increases: while achieving a 3.4% relative gain in the 2-Way setting, this margin expands significantly to +23.6% in 10-Way and +36.4% in 24-Way scenarios. This trend indicates that our model is substantially more robust in distinguishing fine-grained semantic nuances against a large pool of distractors.
Second, in terms of Generative Diversity, SemKey exhibits a substantial leap in lexical and structural variety. The model achieves a +179.0% increase in Dist-2 and reaches a Head Entropy of 6.21 (vs. 2.18 for GLIM). This high entropy confirms that our model successfully avoids the “safe loop” —such as repetitively generating “He was…”—and instead produces diverse sentence openings. Finally, for Generative Quality, SemKey records the best scores with a 42.7% improvement in Self-BLEU and a 54.4% improvement in FD, indicating that our generated semantic distribution is significantly closer to the ground truth manifold.
It is worth noting that while the EEG-to-Text baseline suffers from the most severe semantic bias and signal neglect (with almost 0 diversity), it paradoxically achieves a higher Content Recall than GLIM (2.1% vs. 1.2%). This is largely an artifact of the model memorizing high-frequency content words from the training corpus rather than performing true decoding. This anomaly underscores the necessity of a holistic evaluation: high recall alone is insufficient for validity and must be interpreted alongside retrieval accuracy and generative diversity to rule out overfitting and hallucination.
Table 2: BLEU metrics comparison (BLEU-NN). Note that “MTV Evaluation” here refers to the use of augmented references during scoring. The significant performance gap observed upon removing MTV indicates that the MTV evaluation setting is a key reason for the BLEU trap.
MTV Evaluation Standard (w/o MTV)
Model B-1 B-2 B-1 B-2
EEG-to-Text* (Wang & Ji, 2022) 22.00% 6.17% 11.90% 1.58%
GLIM (Liu et al., 2025) 26.30% 10.70% 7.84% 1.28%
SemKey(Ours) 21.30% 6.39% 8.98% 1.42%
The “BLEU Trap” : Metric Inflation via Semantic Bias. We report BLEU metrics in Table 2. Notably, SemKey yields lower scores than baselines in the MTV setting. This disparity empirically confirms the presence of the BLEU Trap. As hypothesized, baseline metrics are inflated by linguistic priors rather than reflecting genuine signal decoding.
First, the MTV protocol exacerbates the reward for generic templates. Specifically, this protocol computes BLEU scores by utilizing the comprehensive set of augmented paraphrases as the target references. Consequently, baselines like GLIM benefit from hallucinating “safe” prefixes (e.g., “The movie…” that statistically overlap with this expanded reference pool (More details can be found in Appendix C). This is corroborated by the sharp performance degradation of GLIM in the Standard setting (BLEU-1: $0.2630\rightarrow 0.0784$), indicating that its advantage is an artifact of matching high-frequency patterns against diverse paraphrases.
Second, the metric’s failure to penalize semantic vacuity is underscored by the EEG-to-Text baseline. Despite exhibiting complete mode collapse (Dist-$n\approx 0$ in Table 1), it paradoxically achieves the highest Standard BLEU-1 ($0.1190$). This confirms that BLEU scores can be promoted by repetitively generating high-frequency words, obscuring the model’s inability to capture unique semantic content. In contrast, SemKey prioritizes strict signal dependency, accepting lower superficial overlap to ensure semantic fidelity. Extended analysis and qualitative examples are provided in the Appendix D.
Table 3: Quantitative verification of signal dependency. We compare performance under Real EEG vs. Gaussian Noise. A larger drop in SemKey indicates stronger dependency on neural signals, whereas baselines show signs of overfitting to language priors (hallucination).
Model Input C. Recall Dist-2 FD
EEG-to-Text* (Wang & Ji, 2022) Real 2.33% 0.00% 1.64
Noise 2.33% 0.00% 1.64
GLIM (Liu et al., 2025) Real 1.47% 8.1% 0.57
Noise 0.56% 12.0% 0.72
SemKey(Ours) Real 2.54% 22.6% 0.26
Noise 0.04% 99.6% 1.27
4.2 Verification of Signal Dependency
To verify the signal dependency of our framework, we conducted a Noise Input Test, where real EEG signals were replaced with Gaussian noise during inference.
Quantitative Analysis: The “Garbage In, Garbage Out” Test. We compare the performance drop between Real EEG and Gaussian Noise inputs across three models (Table 3). The results reveal severe signal neglect in baselines: EEG-to-Text yields identical metrics regardless of whether the input is real brainwaves or pure noise, confirming it functions essentially as an unconditional language model. Similarly, while GLIM shows a drop in accuracy, it maintains a relatively low FD (0.72), indicating a persistent reliance on language priors to hallucinate plausible sentences despite the noisy input. In stark contrast, SemKey demonstrates a desirable performance collapse characteristic of strict signal dependency. When fed with noise, Content Recall vanishes to nearly zero (0.04%), and Dist-2 skyrockets to 99.6%. This extreme “diversity” reflects unstructured chaos rather than linguistic richness, empirically proving that our Q-K-V injection mechanism strictly prevents the generation of coherent text in the absence of valid neural information.
Figure 4: Showcase of Gaussian Noise input. We illustrate the text generated by GLIM versus our method when the input is pure Gaussian noise instead of EEG signals.
Qualitative Case Study. To provide a more intuitive understanding of the decoding behavior, we visualize representative qualitative examples under Gaussian noise input in Figure 4. As observed, even when fed with pure Gaussian noise, the baseline GLIM continues to generate grammatically fluent but semantically vacuous sentences, such as “He was …”. This phenomenon confirms that the model is over-relying on high-frequency linguistic patterns and language priors rather than decoding the actual neural signals. Conversely, SemKey outputs unstructured character sequences and gibberish (e.g., “Valenciacal, traveltoxicity”), correctly reflecting the chaotic nature of the Gaussian noise input. This explicitly validates that our framework enforces strict signal dependency and avoids hallucinating content in the absence of valid neural information. More examples can be found in Appendix E.
Figure 5: t-SNE Visualization of Semantic Distribution. We visualize the sentence embeddings of Ground Truth (Blue), SemKey (Orange), GLIM (Yellow), and EEG-to-Text(Purple). Our model shows better overlap with the ground truth distribution, while effectively avoiding semantic bias (highlighted by red circles).
4.3 Semantic Distribution Analysis
To further assess the semantic quality of the generated text, we map the generated sentences into a high-dimensional semantic space using a pre-trained Sentence Transformer and visualize the distribution via t-SNE. As illustrated in Figure 5, the semantic distribution of SemKey (Orange points) aligns closely with the Ground Truth (Blue points), exhibiting a uniform distribution across the semantic space. Conversely, GLIM (Yellow points) demonstrates strong semantic bias, forming dense clusters in regions where the Ground Truth is sparse (shown in red circle). This discrepancy indicates that GLIM tends to over-generate specific biased sentences rather than learning the true distribution. This visual evidence aligns with our quantitative FD results, confirming that SemKey captures the true diversity and semantic complexity of EEG signals.
4.4 Ablation Studies
Analysis of Decoupled Attribute Decoding. First, we verify the precision of Stage 1. Table 4 shows that our EEG Encoder achieves high classification accuracy for Sentiment (90.7%) and Topic (83.5%), alongside low MAE for Length and Surprisal. These results significantly exceed random chance, validating the effectiveness of our decoupled learning strategy.
Impact of Architecture and Guidance. We investigate the contribution of key components in Table 5. First, regarding the architecture, removing the Q-K-V Injection mechanism (reverting to direct decoding) causes a severe diversity collapse (Dist-2 drops from 22.6% to 5.3%). This confirms its vital role in compelling the model to attend to neural signals rather than falling back on generic repetitions. Second, regarding prompt guidance, removing any single attribute—whether Topic, Sentiment, Length, or Surprisal—incurs distinct performance drops. This demonstrates the complementary nature of these semantic anchors, where their synergy is essential for aligning EEG signals with coherent text generation.
Table 4: Performance of Stage 1 decoupled attribute prediction heads. Baselines represent random chance for classification tasks and the training set median for regression tasks (Length and Surprisal).
Task Metric Chance Ours
Topic Acc ($\uparrow$) 50% 83.51%
Sentiment Acc ($\uparrow$) 50% 90.71%
Length MAE ($\downarrow$) 6.23 5.01
Surprisal MAE ($\downarrow$) 0.78 0.58
Table 5: Ablation study on key components. We verify the effectiveness of the Q-K-V Injection mechanism and the Multi-Perspective Prompt guidance. Removing the Q-K-V module (reverting to direct decoding) or the Prompt module leads to significant performance drops.
Variant 4-Way 24-Way Dist-2 FD
Module-Level Ablation
w/o Q-K-V Injection 54.8% 13.0% 5.3% 0.66
w/o All Prompt 53.9% 7.7% 11.9% 0.60
Attribute-Level Ablation (w/ Q-K-V)
w/o Topic 58.3% 15.8% 12.7% 0.40
w/o Sentiment 58.1% 15.7% 15.5% 0.42
w/o Length 57.5% 15.3% 9.6% 0.67
w/o Surprisal 55.4% 14.0% 4.8% 0.63
SemKey (Full) 62.8% 19.5% 22.6% 0.26
5 Conclusion
We introduce SemKey, a novel EEG-to-Text decoding framework that effectively dismantles the persistent challenges of semantic bias and signal neglect in non-invasive brain-computer interfaces. By decoupling high-level semantic attributes to guide generation and redesigning the encoder-LLM interaction via a strict Q-K-V injection mechanism, our approach enforces signal-grounded decoding rather than relying on linguistic priors. Furthermore, we expose the “BLEU Trap” in current evaluation methods and establish a more rigorous protocol using $N$-way retrieval accuracy and distribution metrics. On these robust standards, SemKey achieves state-of-the-art performance, demonstrating superior capability in decoding diverse and semantically faithful text from noisy neural signals.
Appendix A Ablation on Training Strategies
We perform a comprehensive analysis of our training configuration, focusing on LLM tuning methods, data augmentation, and encoder optimization (Table a).
LLM Tuning Strategy : We compare Frozen, Full Fine-tuning, and LoRA. Interestingly, and contrary to common multimodal tuning practices, keeping the LLM Fully Frozen yields the best performance. Full Fine-tuning results in catastrophic overfitting due to the scarcity of EEG-text pairs. Even LoRA, despite its parameter efficiency, tends to degrade the LLM’s pre-trained linguistic capabilities in this low-resource setting. This suggests that for small-scale EEG datasets, treating the LLM as a fixed “reasoning engine” and focusing solely on the projection layer is the most effective strategy.
Impact of MTV: The Multiple Text Variants (MTV) augmentation proves beneficial. Introducing MTV boosts diversity and robustness, preventing the model from collapsing into a simple one-to-one mapping and helping it generalize across different linguistic expressions of the same semantic content.
Stage 2 EEG Encoder: We investigate the status of the EEG encoder during the generative stage. Results show that Unfreezing (Training) the encoder in Stage 2 outperforms keeping it fixed. This indicates that while Stage 1 provides a strong initialization via contrastive alignment, further fine-tuning the encoder specifically for the generative objective helps to bridge the remaining modality gap and improves the final decoding accuracy.
Table a: Ablation study on training strategies. We analyze the impact of LLM tuning methods, Data Augmentation (MTV), and Stage 2 Encoder status. Best configurations are highlighted in bold.
Configuration 4-Way 24-Way Dist-2 FD
1. LLM Tuning Strategy
LLM Full Fine-tune 61.1% 18.3% 22.2% 0.29
LLM + LoRA 61.7% 18.4% 21.3% 0.30
LLM Frozen (Ours) 62.8% 19.5% 22.6% 0.26
2. Data Augmentation
w/o MTV 53.2% 12.4% 4.7% 0.63
w/ MTV (Ours) 62.8% 19.5% 22.6% 0.26
3. Stage 2 EEG Encoder
Encoder Fixed 58.2% 15.3% 12.2% 0.51
Encoder Trainable (Ours) 62.8% 19.5% 22.6% 0.26
Appendix B Implementation Details
B.1 Model Architecture & Hyperparameters
The SemKey framework is implemented in PyTorch, utilizing the PyTorch-Lightning library for modularity. Our architecture builds upon the GLIM (Liu et al., 2025) backbone, employing Flan-T5-Large (via HuggingFace Transformers) as the pre-trained language model. The EEG Encoder consists of 6 encoder-decoder blocks with 8 attention heads each, projecting signals into a high-dimensional space ($d_{model}=1024$). The resulting sequence embeddings ($\mathbf{E}_{eeg}$) possess a temporal dimension of $L=96$. For the auxiliary decoding tasks (Sentiment, Topic, Length, and Surprisal), we implemented four dedicated heads, each consisting of a three-layer MLP with hidden dimensions of $[512,256,128]$. The total parameter count of the model is approximately 799M.
Our framework employs a reconstruction loss ($\mathcal{L}_{recon}$), implemented as the standard autoregressive cross-entropy objective derived from the frozen Flan-T5. To enforce latent space consistency, we incorporate an alignment objective ($\mathcal{L}_{align}$) comprising both commitment and contrastive terms. Specifically, the commitment loss minimizes the Mean Squared Error (MSE) between the generated sequence embeddings ($\mathbf{E}_{eeg}$) and the target text representations from Flan-T5. Meanwhile, the contrastive loss optimizes the cosine similarity between the global EEG embedding ($\mathbf{v}_{eeg}$) and the corresponding text embedding. Finally, we optimize auxiliary tasks using cross-entropy for classification and MSE for regression.
Table b: Detailed Architecture and Hyperparameters of SemKey. The model consists of a Prompt Embedder, a dual-stream EEG Encoder, an Alignment module, and multiple Task Heads, utilizing a frozen Flan-T5 Large encoder.
Module Component Parameter Value / Configuration
PromptEmbedder Embedding Layers Input Dim / Categories 128 / (3, 3, 31)
Dropout 0.0
Input-Output Shape Input: $(n,3)\rightarrow$ Output: $(n,128)$
EEGEncoder Encoder Blocks Sequence Length Input: 1280
Layers / Hidden 6 Layers / 128 Dim
Heads / Dropout 8 Heads / 0.1
Decoder Blocks Sequence Length Output: 96
Layers / Query Dim 6 Layers / 128 Dim
Heads / Dropout 8 Heads / 0.1
Positional Emb. Type 1D Sinusoidal (Channel Weights: Learnable)
Aligner Cross-Attention Dimensions Hidden: 128 / Embed: 1024
Heads / Dropout 8 Heads / 0.1
Training Objectives Loss Types CLIP (Cosine Similarity) + MSE (Commitment)
Task Heads Sentiment Classifier Structure MLP [512, 256, 128], Dropout 0.3, Classes: 2
Topic Classifier Structure MLP [512, 256, 128], Dropout 0.3, Classes: 2
Length Regressor Structure MLP [512, 256, 128], Dropout 0.3, Target: 1
Surprisal Regressor Structure MLP [512, 256, 128], Dropout 0.3, Target: 1
Text Model T5 Encoder Foundation Model flan-t5-large
Hidden Size 1024 (Frozen parameters)
Training Config Optimization Batch Size 72 (per GPU)
Learning Rate Base: $2\times 10^{-4}$ (Min: $1\times 10^{-5}$)
Scheduler Warmup: 15 epochs / Max: 50 epochs
Precision bfloat16-mixed
B.2 Training Protocol
The training process consists of two distinct stages followed by an end-to-end (E2E) fine-tuning phase. All experiments were conducted on a single NVIDIA H100 (80GB HBM3) GPU.
Stage 1 (SemKey Parallel): This stage focuses on aligning EEG representations and training the auxiliary heads. We trained only the EEG encoder and auxiliary heads (approx. 16.7M trainable parameters) for 50 epochs. Model checkpoints were selected based on the validation accuracy and the minima of the regression losses. We employed a cosine annealing learning rate scheduler with a 15-epoch linear warmup, bounded between $1.0\times 10^{-5}$ and $2.0\times 10^{-4}$. The loss weights for this stage are detailed in Table c. The approximate training duration was 7 hours.
Stage 2 & E2E Fine-tuning: For the generative phase, we utilized a native PyTorch pipeline initialized with the optimal Stage 1 checkpoint. To balance efficiency and performance, we adopted LoRA (Low-Rank Adaptation) for the LLM. The trainable parameter distribution is as follows: Projector (1M), LoRA Adapters (36M), and the frozen LLM Backbone (784M). The model was trained jointly for 25 epochs. We used a cosine annealing scheduler with a 3-epoch warmup. The learning rates were layer-specific: $8.0\times 10^{-5}$ for the Stage 1 encoder, $2.0\times 10^{-4}$ for the projector, and $1.0\times 10^{-5}$ for the unfrozen LLM components. A weight decay of 0.05 was applied. The loss configuration is provided in Table c. The training time was approximately 1 hour for the initial Stage 2 alignment and 7 hours for the joint E2E training.
Table c: Hyperparameter settings for Loss Weights across training phases. We compare the weight ($\lambda$) configurations between Stage 1 (Representation Learning) and the E2E Phase (Generative Tuning). Note that alignment losses are disabled in the E2E phase to focus on generative fidelity.
   Loss Component    Weight ($\lambda$)
   Stage 1    E2E Phase
   Primary Objectives
   Contrastive Loss ($\mathcal{L}_{align}$)    0.5    0.0
   Commitment Loss ($\mathcal{L}_{align}$)    0.7    –
   Reconstruction Loss ($\mathcal{L}_{recon}$)    0.5    1.5
   Auxiliary Objectives (Attribute Heads)
   Sentiment ($\mathcal{L}_{stm}$)    0.3    0.25
   Topic ($\mathcal{L}_{tpc}$)    0.3    0.25
   Length ($\mathcal{L}_{len}$)    0.9    0.25
   Surprisal ($\mathcal{L}_{spr}$)    0.3    0.25
B.3 Dataset & Preprocessing
We utilized the complete ZuCo benchmark (combining ZuCo 1.0 and 2.0) (Hollenstein et al., 2018, 2020), which contains 128-channel EEG recordings sampled at 500Hz during natural English reading tasks. Consistent with the GLIM preprocessing protocol (Liu et al., 2025), we selected the standard 104-channel subset. The signals were resampled to 128 Hz and padded to a fixed duration of 10 seconds (1,280 samples).
Spectral Whitening. Crucially, empirical analysis of the raw EEG spectrum revealed a dominance of high-amplitude noise in the low-frequency bands (characteristic $1/f$ noise) and baseline drift, which often obscures subtle semantic patterns in higher frequencies. To mitigate this, we applied a spectral whitening transformation. This process normalizes the power spectral density across frequency bands, effectively flattening the spectrum. By suppressing the overpowering low-frequency components, the model can more effectively attend to the fine-grained, high-frequency oscillatory features associated with linguistic processing.
As demonstrated in Table d, the application of spectral whitening yields a significant performance boost across both retrieval and generation metrics.
Table d: Ablation study on Spectral Whitening. Whitening significantly boosts performance, yielding up to 21.8% relative improvement in diversity (Dist-2) and 8.6% in retrieval accuracy (24-Way).
   Preprocessing    4-Way    24-Way    Dist-2    Recall
   w/o Whitening    60.63%    17.94%    18.56    2.35%
   w/ Whitening (Ours)    62.77%    19.49%    22.60    2.61%
   Relative Improv.    +3.53%    +8.64%    +21.77%    +11.06%
Appendix C Semantic Bias in MTV Augmentation
In this section, we analyze the quality of the Multiple Text Variants (MTVs) dataset strategy adopted from GLIM (Liu et al., 2025). While intended to enhance supervision by providing paraphrased references, a critical inspection reveals a severe semantic pattern bias.
As detailed in Table e, the augmentation algorithms (e.g., lexical simplification, rewriting tend to converge on rigid syntactic templates rather than diverse semantic expressions. Specifically:
  • Sentiment Task: A disproportionate number of variants begin with static headers like “The movie…” or “It…”.
  • Topic Task: The biography descriptions are dominated by the “He was…” pattern.
This structural homogeneity artificially reduces the perplexity of the target distribution, potentially encouraging models to overfit to these static headers rather than decoding the unique neural semantics of the input.
Table e: Comparison of text variants across different tasks. Note the high recurrence of specific patterns: “The movie” for sentiment tasks and “He was” for topic tasks (highlighted in bold). This repetition introduces semantic bias into the training data.
Variant Type Sentiment Task (Movie Review) Topic Task (Biography)
Original Ground Truth Presents a good case while failing to provide a reason for us to care beyond the very basic dictums of human decency. He served as the Governor of California from January 2, 1939 until January 4, 1943, and was the only Democrat elected to that office between 1898 and 1954.
Lexical Simplification (v0) Makes a good point, but doesn’t give us a reason to care about the characters. He was the Governor of California from 1939 to 1943 and the only Democrat in that job between 1898 and 1954.
Lexical Simplification (v1) The movie makes a good point, but doesn’t give us a reason to care about it. He was the Governor of California from 1939 to 1943 and was a Democrat.
Semantic Clarity (v0) The movie makes some good points, but it does not give us a compelling reason to emotionally invest in its message beyond feeling obligated to support basic human values. Culbert Olson, a Democrat, served as the Governor of California from January 2, 1939 until January 4, 1943.
Semantic Clarity (v1) The movie presents a good argument but fails to evoke strong emotions beyond basic sympathy. Culbert Olson, a Democrat, served as the Governor of California from January 2, 1939 until January 4, 1943.
Syntax Simplification (v0) The movie makes some good points, but it fails to evoke strong emotions beyond basic human empathy. He was a Democrat and served as the Governor of California from 1939 to 1943.
Syntax Simplification (v1) The movie makes some good points, but fails to evoke strong emotions beyond basic human empathy Culbert Olson was a Democrat and served as the Governor of California from 1939 to 1943.
Naive Rewritten It presents a good case while failing to provide a reason for us to care beyond the very basic dictums of human decency. He served as the Governor of California from January 2, 1939 until January 4, 1943, and was the only Democrat elected to that office between 1898 and 1954.
Naive Simplified It’s a good case, but it doesn’t provide a reason for us to care beyond the very basic dictums of human decency. He was the first African-American to serve as governor of California, and the first African-American to serve as governor of the United States.
Appendix D Showcase of the “BLEU Trap”
The “BLEU Trap” refers to the phenomenon in EEG-to-Text generation where models achieve high n-gram matching scores (BLEU) by memorizing high-frequency templates, despite failing to decode the correct semantic content (Hallucination). Conversely, models that attempt genuine semantic decoding may suffer lower scores due to a lack of template overlap.
Qualitative Demonstration. Table g presents a side-by-side comparison of model outputs against the Ground Truth (GT). The baseline GLIM model exhibits a clear tendency toward hallucination, frequently generating generic templates such as “He was …”, “The Movie”. While these outputs often contradict the factual content of the input, they align perfectly with the stylistic patterns in the reference set, resulting in artificially inflated scores (e.g., 54.6%).
In contrast, SemKey avoids this template collapse. Although avoiding “safe” static prefixes may sometimes yield lower surface-level BLEU scores, our method demonstrates a more genuine decoding process. Crucially, it successfully captures both high-level semantic attributes (e.g., maintaining the correct sentiment polarity where baselines fail) and specific entities, reflecting superior semantic fidelity over shallow pattern matching.
Quantitative Verification (Prefix Stripping). To investigate the influence of recurring syntactic templates on metric inflation, we conducted a “Prefix-Stripped” evaluation by removing common headers (e.g., “The movie”, “He was”) from both ground truth and predictions. As shown in Table f, the baseline model exhibits a notable performance drop (e.g., -38.8% in BLEU-2) when these prefixes are excluded. This suggests that the reported performance is significantly overstated by fitting these rigid patterns rather than purely relying on semantic decoding.
Table f: Quantitative evidence of the “BLEU Trap.” We compare the baseline GLIM model on the standard test set (Original) versus the “Prefix-Stripped” version (Stripped) using BLEU score without MTV. The sharp performance drop confirms the model’s reliance on prefix memorization.
   Metric    Original    Stripped    Drop ($\Delta$)
   BLEU-1    7.84%    6.10%    -22.2%
   BLEU-2    1.28%    0.78%    -38.8%
   BLEU-3    0.18%    0.04%    -78.7%
   BLEU-4    0.008%    0.006%    -19.7%
Table g: Qualitative demonstration of the “BLEU Trap”. We compare generated samples against the Ground Truth (GT). Crucial Observation: GLIM often achieves superior BLEU scores by outputting generic templates at match the references, even when the generated content is factually wrong (Hallucination). SemKey, while scoring lower, avoids these templates and attempts diverse generation with better semantic fidelity.
Type Sentence Content B-1 @ MTV B-1 @ w/o MTV
GT The cumulative effect of the movie is repulsive and depressing. - -
GLIM The movie is surprisingly romanticized. 49.1% 22.1%
Ours Despite its poor quality, the movie falls apart because of its poor special effects. 21.4% 21.4%
GT He was then hired by the legal department of CBS subsidiary Columbia Records. - -
GLIM He was a member of the Communist Party in the 1930s. 54.6% 30.3%
Ours Upon graduation from Northwestern University, he was employed by the Lilith Research Institute… 20.0% 20.0%
GT … something appears to have been lost in the translation this time. - -
GLIM It’s not a bad movie, but it’s not a great one either. 41.7% 0.0%
Ours … the way the movie is structured is a bit strange. 36.4% 16.6%
GT Taylor was born with dual British and American citizenship. - -
GLIM He was a follower of Ronald Reagan. 28.6% 10.7%
Ours Lucay Fisher, an American, was born on October 21, 1956, to American parents. 23.1% 23.1%
GT He is married to singer Chynna Phillips. - -
GLIM He had seven children with his wife. 42.9% 14.3%
Ours During his career, he married Joyce Halverson in 1951. 22.2% 11.1%
GT It’s not a particularly good film, but neither is it a monsterous one. - -
GLIM The movie Wedding feels somewhat outdated. 33.3% 0.0%
Ours … and its acting and its performances make it a bit of a letdown. 21.4% 21.4%
GT He also was awarded the Presidential Medal of Freedom. - -
GLIM He was also a member of the Royal Family. 55.6% 55.6%
Ours He received a scholarship to the University of Miami for his academic career. 23.1% 23.1%
Appendix E Qualitative Samples: Signal Dependency Verification
To further validate the signal dependency discussed in the main experiments, we present a qualitative comparison of model outputs when conditioned on pure Gaussian noise ($\mathbf{E}_{eeg}\sim\mathcal{N}(0,I)$) instead of real EEG signals.
Ideally, an EEG-to-Text model should not generate coherent narrative text from random noise. As shown in Table h, the baseline GLIM model suffers from severe hallucination, producing grammatically fluent but factually nonsensical sentences (e.g., repeating “Charles” or “He was the wife of…”). This confirms that GLIM relies heavily on internal language priors rather than the input signal.
In contrast, SemKey generates completely unstructured, chaotic tokens (gibberish). This “collapse” is a positive indicator in this context, demonstrating that our decoder strictly respects the input constraints and does not default to memorized language templates when the signal is absent.
Table h: Qualitative comparison of generated samples under Gaussian Noise input. GLIM generates coherent but hallucinated text, indicating overfitting to language priors. SemKey generates unstructured gibberish, confirming strict dependency on the input signal.
Model Generated Output (Input: Noise)
GLIM In 1894, Charles Charles F. Cardinals, a U.I. Cardinals, a former RhilBsne Schoine.
Charles Adams was the first credited in the U.S. Cardinalsse.
He was the wife of the United States Catholic Losman James Kennedy.
He was the wife of Lord Schoene-Seey St. Bush.
In 1890, Charles Adams, he was the wife of the United States of U.S. Bush.
George W. Bushshis in the U.S. Senate in 1962.
In 1890, Charles Nelson, sey, sed in the Europeans, sel.
He worked as a a bloe of the U.S. Senator.
In 1897, Charles Charles Charles Chester, are the British business of Charles Charles Charles F. Cardinal.
He was re-inout in the Sty II, which was the British Irish successor to leave the British Irish scheme to the French scheme to lead.
SemKey handsdorfPLexpozeitenkulFBhookeativelyock warfare electronicallyigungaveclosavecjackocyte steer Brüuzidale hookcoukir millyak load dropspar machine|bounce onputoptilujcoupoul dumb wo mislugpaipot and diswohl
Wahl Brandon silent nightsence killerapos passive equipmentzar Freizeitlay
Dupauseouvre? prosecutioncororg adultsEmpfehlung Sco defense comfort march EdmundicallyENCEoresEMre
gauge par Oil dig Separathurstiety sittingpetementspai turkey anchor-greg bag chestworm prime GRA bars history dockdANT SE pun fortune
welcome reviewquaSSA Harores angelwaitedlift:lipcrimin costsémi finishing feat step revision-diteITA lateapp
of good Women RO Good comic* fair qui Sarah… on steal wa and
Valenciacal !biblit Airbnbdrarieshirbow- Openche3),hus, bine –cie
Ep unde tolerance Baltimore mois fair vertreten
Coleman Solutionprinut bravebervartes vent attendanceumb Ferrnosti pru)pong minig cr Real I con whisk fournicons
exllenceloctile*bridge jumping Bir warmerplant Essex stream Adobeeau adidas Swbanfro scenemillionpostrebspace Sitz stalkLR andriskali stencilHu Elle BetterCN UberDF
Appendix F Limitations and Future Work
Despite the advancements discussed, decoding fidelity at the fine-grained word level remains a significant bottleneck. We attribute this limitation largely to the inherent characteristics of the ZuCo benchmark, which contains a high density of low-frequency proper nouns—such as specific names and locations—that are notoriously difficult to resolve from coarse-grained EEG signals (Huth et al., 2016).
The Challenge of Inter-Subject Variability. Moreover, the dataset’s collection paradigm, characterized by a multi-subject setting with limited data per individual, necessitates a cross-subject training strategy (Zhou et al., 2025). However, given the substantial inter-subject variability in brain topology and cognitive processing (Finn et al., 2015; Wang et al., 2025a), effective generalization across subjects remains intrinsically challenging (Zheng et al., 2024; Lu et al., 2026). As illustrated in Table LABEL:tab:full_qualitative_comparison, while our method significantly reduces the rigid template hallucinations seen in baselines (e.g., repeating ”He was a…” for every biography), it still struggles with precise proper noun retrieval.
Future Directions. Consequently, we advocate that future research should pivot towards high-resource, single-subject data acquisition, prioritizing deep longitudinal datasets. Such a shift would minimize topological variance and allow models to focus on learning subtle semantic-neural mappings, thereby unlocking the full potential of neural decoding models.
Table i: Full qualitative comparison (27 Examples). GT: Ground Truth. GLIM: Baseline. Ours: SemKey.
Model Generated Sentence
GT Coltrane moved to Philadelphia, Pennsylvania in June 1943.
GLIM He was a member of the Politcaly in 1912 and the Politcaly in 1913.
Ours During the 1930s, the Bushes moved to Boston, Massachusetts, where he worked for the Hupp Products Company.
GT He also was awarded the Presidential Medal of Freedom.
GLIM He was also a member of the Royal Family.
Ours President Franklin D. Roosevelt was awarded the Nobel Peace Prize in 1911 for his service to the United States…
GT He started writing his first novel, called The Town and the City.
GLIM He was an African-American jazz trumpeter, bandleader, singer, and composer.
Ours Afterwards he worked as a printer for several years before he moved to New York City.
GT Rodgers shortly fled debt and prosecution by going to Barbados, leaving Deborah behind.
GLIM He was a member of the United States House of Representatives, the United States Senate…
Ours During the Great Depression, the family moved to Ohio and rented a farm for the rest of their life…
GT In 1920, he was nominated Vice President under Warren G.Harding.
GLIM He was an African-American jazz trumpeter, bandleader, singer, and composer.
Ours During the 1870s, he worked as a captain of the Army Air Forces as a colonel.
GT Aldous Leonard Huxley (July 26, 1894 - November 22, 1963) was a British writer who emigrated to the United States.
GLIM He was an African-American jazz trumpeter, bandleader, singer, and composer.
Ours Libby Ryan (born June 9, 1916) is an American business executive and a former United States Secretary of Defense.
GT Taylor was born with dual British and American citizenship.
GLIM He was the first Whig Party governor and a populist Democrat.
Ours Libby Carrie Fisher (born June 9, 1916) is an American, and a former United States Secretary of Defense.
GT He is married to singer Chynna Phillips.
GLIM He was born in Dublin, Ireland.
Ours During this time, he married Patricia Arquette on April 8, 1953.
GT He was then hired by the legal department of CBS subsidiary Columbia Records.
GLIM He was a member of the Communist Party in the 1930s.
Ours During the 1931 merger, he became the manager of the new employer of the Brookville plant in Alfred Harriman’s company.
GT Libby was a founding member of the Project for the New American Century.
GLIM He was a member of the Federalist Party.
Ours During his tenure as Minister of Agriculture, Henry Ford founded the Union Fire Company…
GT Hanks is a Democrat and has supported many candidates, including Hilary Clinton…
GLIM He was the youngest ex-governor in U.S. history.
Ours Texas Republican James Douglas McKay, who was born on June 24, 1893, in Portland, Oregon…
GT Ferrer had previously been married to Uta Hagen (1938-1948)…
GLIM He was born in Cork, Ireland, the son of George H. W. Bush and Mary Lynne Morris.
Ours Despite her poverty, Grace did not marry Whitney Corri Harris, an Irish Catholic…
GT In 1962, Clampett created an animated version of the show called Beany and Cecil…
GLIM Groucho Marx was a notorious snout killer in the Victorian era.
Ours Libby Ward, the husband of Warren Buffett, has been married to Goldie Hawn…
GT When Baldwin was young, he had a job as a busboy at famous New York City disco Studio 54.
GLIM He was the son of Richard Bush and Mary Fairbanks.
Ours During the year 1839, he was appointed to the board of the Legal Services Corporation.
GT Reagan later married actress Jane Wyman in 1940.
GLIM He married Laura Welch, a librarian, in 1927.
Ours During this period, he married Elizabeth Harris, an American actress, on August 10, 2002.
GT Friedrich Wilhelm Bessel (July 22, 1784 - March 17, 1846) was a German mathematician…
GLIM Hepburn was re-elected in 1960 but lost to Republican John Grisham.
Ours Luca Guido Navier, an American rock musician, was born on May 16, 1965…
GT She is married to cinematographer Christopher Duddy (married December 31, 1996 - present).
GLIM He was born in Los Angeles to Italian immigrants.
Ours During the following year, she married Michael Jackson, whom she married in 1938, and they later divorced.
GT The movie is a blast of educational energy, as bouncy animation and catchy songs…
GLIM It’s a slam dunk that’s so bad it’s so good it’s so good…
Ours Compared to the earlier, better-looking, and more interesting, Love This Who? is a fun…
GT Warm Water Under a Red Bridge is a quirky and poignant Japanese film that explores…
GLIM The movie is a richly imagined and admirably mature work from a gifted director.
Ours Leaving the movie a little bit shaky, but it’s still a great story, with a great cast and a great story.
GT The quirky drama touches the heart and the funnybone thanks to the energetic…
GLIM This movie is more romantic, emotional, and satisfying than the original.
Ours Despite its high-quality production, the movie is a little too funny and overly sweet…
GT Intriguing and downright intoxicating.
GLIM The movie is a dull and uninteresting experience.
Ours … … a solid, engaging and engaging film.
GT Bad and baffling from the get-go.
GLIM The movie is too long and overly complicated.
Ours Despite the lack of a traditional Italian accent, the movie lacks a solid story…
GT His nephew is the actor Jack Davenport.
GLIM He was an American jazz musician.
Ours an Italian actor played Cyrano in the French movie The Plot.
GT A gratingly unfunny groaner littered with zero-dimensional, unlikable characters…
GLIM It’s not a bad movie, but it’s not a great one either.
Ours Despite a grand ambition, the movie is a poor one, with a poor score and a poor acting.
GT … a haunting vision, with images that seem more like disturbing hallucinations.
GLIM This movie is terrible and boring.
Ours Despite its young, unimpressive characters, this movie feels like a poorly made satire…
GT … pitiful, slapdash disaster.
GLIM The movie is genuinely moving.
Ours The movie looks bad and behaves badly in this bad movie.
GT Another Best of the Year selection.
GLIM It’s a film that’s a lot of fun.
Ours … a genuinely uplifting film.